I need help with these two, they're not in the textbook (GEOMETRY) they're on a review sheet I have to do to prepare for our test tomorrow.
#1) The measure of the supplement of an angle is 32 less than the measure of the angle. Find the measures of the angles.
#2)The measure of one angle is 12 degrees greater than the measure of its complement. What are the measures of the angles?
PLEASE HELP! Thanks!

Answer by Earlsdon(6098)

#1) Let the angle be A, then:
180-A = A-32 "...the supplement of an angle (180-A) is 32 degrees less than the measure of the angle."
180 = 2A-32
212 = 2A
A = 106 degrees and its supplement, 180-A = 74 degrees.
#2) Let the angle be B, then:
B = (90-B)+12 "The measure of one angle (B) is 12 degrees greater than the measure of its complement (90-B)."
B = 78-B Add B to both sides.
2B = 78 Divide both sides by 2.
B = 39 and its complement, 90-B = 51 degrees.